Kofi and Ama have 120 hours each and they can spend their time planting Rice or Maize.
You are told that the opportunity cost of producing one bag of rice for Ama is 1/3 bag of
maize and the opportunity cost of producing one bag of rice for Kofi is ¼ bag of maize.
Ama can produce 6 bags of maize when she uses all her resource for maize production
and 1Kofi can produce 8 bags of rice when he uses all his resource .Use this to answer
the following questions:
a) If Ama decides to use all her resources for rice production, how many bags of rice will she
produce?
b) If Kofi decides to use all his resources for maize production how many bags of rice will he
produce
c) Who has the comparative advantage in planting rice?
d) Who has the comparative advantage in planting maize?
e) What is Kofi’s opportunity cost for planting maize?
f) What is Ama’s opportunity cost for planting TWO bags of rice?
g) Who has Absolute Advantage (AA) in Maize and who has AA in Rice? �
1
Expert's answer
2018-09-26T09:49:08-0400
a) If Ama decides to use all her resources for rice production, then she will produce 6/(1/3) = 18 bags of rice.
b) If Kofi decides to use all his resources for maize production, then he will produce 0 bags of rice.
c) Kofi has the comparative advantage in planting rice.
d) Ama has the comparative advantage in planting maize.
e) Kofi’s opportunity cost for planting maize is 4 bags of rice.
f) Ama’s opportunity cost for planting TWO bags of rice is 2*1/3 = 2/3 bags of maize.
g) Ama has Absolute Advantage (AA) in Maize and Rice, because she can produce either 18 bags of rice or 6 bags of maize, but Kofi can only produce 8 bags of rice or 2 bags of maize.